grade B syrup sells for $1.35/L less then grade A syrup.if 80L of grade B syrup is mixed with 40 L of Grade A syrup, the resulting mixture will be worth $1.85/L. what is the unit price for the two grades of syrup?
Answer by ankor@dixie-net.com(22740)

grade B syrup sells for $1.35/L less then grade A syrup.if 80L of grade B syrup is mixed with 40 L of Grade A syrup, the resulting mixture will be worth $1.85/L. what is the unit price for the two grades of syrup?
:
Let x = price of grade A syrup
Then (x-1.35) = price of the grade B syrup
:
B stuff + A stuff = $1.85 stuff
80(x-1.35) + 40x = 1.85(80+40)
:
80x - 108 + 40x = 1.85(120)
:
120x - 108 = 222
:
120x = 222 + 108
:
120x = 330
:
x = 330/120
:
x = $2.75, price of the A syrup
:
2.75 - 1.35 = $1.40 is the price of B syrup
:
:
Check:
1.4(80) + 2.75(40) = 1.85(120)
112 + 110 = 222; confirms our solutions.